OpenStudy (anonymous):

CORRECTION: what is wrong with solving the equation X^2 = 3x by dividing both sides of the equation by X? Explain the correct method for solving this equation.

OpenStudy (anonymous):

Yep thats wrong.

OpenStudy (amistre64):

as long as x not equal 0, it should be fine. but you lose a solution in the process

OpenStudy (anonymous):

The correct METHOD is: x^2=3x x^2-3x=0 x(x-3)=0 Either x=0 OR x-3=0 Thus. x=0 or x=3

OpenStudy (anonymous):

got it? @ANAID

OpenStudy (anonymous):

Thanks, we came up with the same answer.

OpenStudy (anonymous):

Welcome